# Canonical Form and Finite Blocklength Bounds for Stabilizer Codes

### v1

@@ 5 Finite blocklength bounds for stabilizer codes and Pauli noise
The preparatory subsection 5.1 establishes a property of the uniform distribution over the symplectic group that is used to derive the achievability bound. Another preparatory subsection 5.2 establishes notation necessary to state the bounds. Next, subsection 5.3 states and proves achievability and converse bounds on $\varepsilon_{errorguess}\left(p_{UV},r\right)$ and $R_{errorguess}\left(p_{UV},\epsilon\right)$ for arbitrary $p_{UV}$. Subsection 5.4 and subsection 5.5 show how the general bounds can be computed efficiently in the cases of the qubit erasure channel and the qubit depolarizing channel respectively. Moreover, it is seen there that the achievability and converse bounds on the rate differ by $O(1/n)$ for $n$ qubits.
@@ 5.1 The uniform distribution over symplectic matrices
The proof of Lemma 12 uses the fact that $\mathcal{SP}_{2n}$ acts transitively on $\mathbb{F}_{2}^{2n}\backslash\{0\}$. Previous expositions of the hashing bound also implicitly establish and use this transitive action; see for example [3, Section 3.2.3]. A short proof based on the symplectic Gaussian moves is as follows:
@@ 5.3 General bounds
where $m=n-k$ and $\mathbb{I}\left(\cdot\right)$ is the indicator of an event. Then,
@@ 5.4 The qubit erasure channel
$\varepsilon_{errorguess}^{converse}\left(p_{UV},r\right)=\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)\\ -2^{m}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)$ (105)
@@
$\varepsilon_{errorguess}^{achievability}\left(p_{UV},r\right)=\left(1+\frac{1}{2^{m+1}}\right)\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)-\frac{1}{2^{m+1}}\\ -\frac{2^{m}+1}{2}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)\\ +\frac{(1+3\delta)^{n}}{2^{m+1}}\mathcal{F}\left(n,\frac{4\delta}{1+3\delta},\left\lfloor\frac{m}{2}\right\rfloor\right)$ (106)
where $r=k/n$ and $m=n-k$.
@@
[…]
@@
$-\mathbb{E}\left(\mathbb{I}\left(2|V|>m\right)2^{m-2|V|}\right)\\ =-2^{m}\sum_{i=0}^{n}{n\choose i}\delta^{i}4^{-i}(1-\delta)^{n-i}\mathbb{I}\left(2i>m\right)\\ =-2^{m}\left(\frac{4-3\delta}{4}\right)^{n}\sum_{i=0}^{n}{n\choose i}\left(\frac{\delta}{4-3\delta}\right)^{i}\left(\frac{4-4\delta}{4-3\delta}\right)^{n-i}\mathbb{I}\left(2i>m\right)\\ =-2^{m}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)$ (111)
@@
$\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)(J-1)2^{-m}\right)\\ =\sum_{t}\mathbb{P}\left(V=v\right)\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)(J-1)2^{-m}\Big{|}V=v\right)$ (112)
@@
$\varepsilon_{errorguess}^{achievability}\left(p_{UV},r\right)=\mathbb{E}\left(\mathbb{I}\left(2|V|>m\right)\left(1-2^{m-2|V|-1}-2^{-2|V|-1}\right)\right)\\ +\mathbb{E}\left(\mathbb{I}\left(2|V|\leq m\right)\left(2^{2|V|-m-1}-2^{-m-1}\right)\right)$ (114)
@@
$\mathbb{E}\left(\mathbb{I}\left(2|V|>m\right)\left(1-2^{m-2|V|-1}-2^{-2|V|-1}\right)\right)\\ =\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)\\ -\frac{2^{m}+1}{2}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)$ (115)
@@
$-2^{-m-1}\mathbb{E}\left(\mathbb{I}\left(2|V|\leq m\right)\right)=-2^{-m-1}\mathcal{F}\left(n,\delta,\left\lfloor\frac{m}{2}\right\rfloor\right)\\ =-2^{-m-1}\left(1-\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)\right)$ (117)
@@ 5.5 The qubit depolarizing channel
For every $n,p$, extend the binomial CDF $\mathcal{F}\left(n,p,\cdot\right)$ from equation (104) to a piecewise linear function whose graph connects the points $(-1,0)$, $(0,\mathcal{F}\left(n,p,0\right))$, …, $(n,\mathcal{F}\left(n,p,n\right))$. This extension is an increasing bijection that maps $[-1,n]$ to $[0,1]$; let $\mathcal{F}^{-1}\left(n,p,\cdot\right)$ be the inverse function.
@@
For each $i=0$, …, $n$, the events $J/4^{n}\leq\mathcal{F}\left(n,3/4,i\right)$ and $|U|\leq i$ coincide. Therefore, the CDF of $J/4^{n}$ can be extended to a piecewise linear function $[0,1]\rightarrow[0,1]$ whose graph connects the points $(0,0)$, $(\mathcal{F}\left(n,3/4,0\right),\mathcal{F}\left(n,\delta,0\right))$, …, $(\mathcal{F}\left(n,3/4,n\right),\mathcal{F}\left(n,\delta,n\right))$. This can equivalently be expressed as the composition of two piecewise linear functions: $\mathcal{F}^{-1}\left(n,3/4,\cdot\right)$, whose graph connects the points $(0,-1)$, $(\mathcal{F}\left(n,3/4,0\right),0)$, …, $(\mathcal{F}\left(n,3/4,n\right),n)$, and $\mathcal{F}\left(n,\delta,\cdot\right)$, whose graph connects the points $(-1,0)$, $(0,\mathcal{F}\left(n,\delta,0\right))$, …, $(n,\mathcal{F}\left(n,\delta,n\right))$. ∎
@@
where $r=k/n$, $m=n-k$ and $\ell=\mathcal{F}^{-1}\left(n,\frac{3}{4},\frac{2^{m}}{4^{n}}\right)$.
@@
$R_{errorguess}\left(p_{U},\epsilon\right)=1-h(\delta)-\delta\log_{2}(3)\\ -\sqrt{\frac{\delta(1-\delta)}{n}}\Phi^{-1}\left(\epsilon\right)\log_{2}\left(\frac{\delta}{3(1-\delta)}\right)+\frac{\log_{2}(n)}{2n}+O\left(\frac{1}{n}\right)$ (133)
@@
$\mathbb{P}\left(J>2^{m}\right)+\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)(J-1)2^{-m}\right)\\ =\mathbb{P}\left(J>2^{m}\right)-\frac{1}{2^{m+1}}\mathbb{P}\left(J\leq 2^{m}\right)+\frac{1}{2^{m}}\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)\left(J-\frac{1}{2}\right)\right)\\ =\left(1+\frac{1}{2^{m+1}}\right)\mathcal{F}\left(n,1-\delta,n-1-\ell\right)-\frac{1}{2^{m+1}}\\ +\frac{1}{2^{m}}\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)\left(J-\frac{1}{2}\right)\right)$ (135)
@@
$\mathbb{E}\left(\mathbb{I}\left(4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)<J\leq 4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)\right)\left(J-\frac{1}{2}\right)\right)\\ =\mathbb{P}\left(4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)<J\leq 4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)\right)\\ *\mathbb{E}\left(J-\frac{1}{2}\Bigg{|}4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)<J\leq 4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)\right)\\ =\left(4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)-4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)\right)\left(\frac{\delta}{3}\right)^{i}(1-\delta)^{n-i}\\ *\frac{4^{n}\mathcal{F}\left(n,3/4,i-1\right)+4^{n}\mathcal{F}\left(n,3/4,i\right)}{2}\\ =\frac{(16-16\delta)^{n}}{2}\left(\frac{\delta}{3-3\delta}\right)^{i}\left(\mathcal{F}\left(n,\frac{3}{4},i\right)^{2}-\mathcal{F}\left(n,\frac{3}{4},i-1\right)^{2}\right)$ (138)
@@
$\frac{1}{2^{m}}\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)\left(J-\frac{1}{2}\right)\right)=\frac{(16-16\delta)^{n}}{2^{m+1}}\left(\frac{\delta}{3-3\delta}\right)^{\left\lfloor\ell\right\rfloor+1}\mathcal{F}\left(n,\frac{3}{4},\ell\right)^{2}\\ +\frac{(16-16\delta)^{n}}{2^{m+1}}\left(\frac{3-4\delta}{3-3\delta}\right)\sum_{i=0}^{\left\lfloor\ell\right\rfloor}\left(\frac{\delta}{3-3\delta}\right)^{i}\mathcal{F}\left(n,\frac{3}{4},i\right)^{2}$ (140)
@@
(${n\choose\left\lfloor\ell\right\rfloor}\left(\frac{3}{4}\right)^{\left\lfloor\ell\right\rfloor}\left(\frac{1}{4}\right)^{n-\left\lfloor\ell\right\rfloor}\leq\mathcal{F}\left(n,\frac{3}{4},\ell\right)\leq\frac{3}{2}{n\choose\left\lceil\ell\right\rceil}\left(\frac{3}{4}\right)^{\left\lceil\ell\right\rceil}\left(\frac{1}{4}\right)^{n-\left\lceil\ell\right\rceil}$ (151)
@@
and the ratio of the $(i-1)$-st to the $i$-th term in the sum is
($\frac{{n\choose i-1}\left(\frac{3}{4}\right)^{i-1}\left(\frac{1}{4}\right)^{n-i+1}}{{n\choose i}\left(\frac{3}{4}\right)^{i}\left(\frac{1}{4}\right)^{n-i}}=\frac{i}{3(n-i+1)}\leq\frac{2\delta}{3(1-2\delta)}\leq\frac{1}{3}$ (153)
@@
$\log_{2}\mathcal{F}\left(n,\frac{3}{4},\ell\right)=-n\mathbf{D}\left(\frac{\ell}{n}\middle\|\frac{3}{4}\right)-\frac{1}{2}\log_{2}(n)+O(1)$ (154)
where $\mathbf{D}\left(\cdot\middle\|\cdot\right)$ is the relative entropy.
@@
$\mathbf{D}\left(\frac{\ell}{n}\middle\|\frac{3}{4}\right)=\mathbf{D}\left(\delta\middle\|\frac{3}{4}\right)-\sqrt{\frac{\delta(1-\delta)}{n}}\Phi^{-1}\left(\epsilon\right)\log_{2}\left(\frac{\delta}{3(1-\delta)}\right)+O\left(\frac{1}{n}\right)$ (156)
@@
$\mathbf{D}\left(x\middle\|y\right)=\mathbf{D}\left(x_{0}\middle\|y\right)+(x-x_{0})\log_{2}\left(\frac{x_{0}(1-y)}{(1-x_{0})y}\right)+O\left((x-x_{0})^{2}\right)$ (157)
@@
Claim 4: (154) and (156) and the identity $2-\mathbf{D}\left(\delta\middle\|3/4\right)=h(\delta)+\delta\log_{2}(3)$ prove (149).

### v2

@@ 5 Finite blocklength bounds for stabilizer codes and Pauli noise
The preparatory subsection 5.1 establishes a property of the uniform distribution over the symplectic group that is used to derive the achievability bound. Another preparatory subsection 5.2 establishes notation necessary to state the bounds. Next, subsection 5.3 states and proves achievability and converse bounds on $\varepsilon_{errorguess}\left(p_{UV},r\right)$ and $R_{errorguess}\left(p_{UV},\epsilon\right)$ for arbitrary $p_{UV}$. Subsection 5.4 and subsection 5.5 show how the general bounds can be computed efficiently in the cases of the qubit erasure and depolarizing channels respectively. Moreover, it is seen there that the achievability and converse bounds on the rate differ by $O(1/n)$ for $n$ qubits. Subsection 5.6 shows that the achievability and converse bounds can be computed in polynomial time in the case of $n$ independent identical Pauli channels with side information.
@@ 5.1 The uniform distribution over symplectic matrices
[…]
@@ 5.3 General bounds
where m=nkm=n-k and $\mathbb{I}\left(\cdot\right)$ is the indicator of an event. Then,
@@ 5.4 The qubit erasure channel
$\varepsilon_{errorguess}^{converse}\left(p_{UV},r\right)\\ =\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)-2^{m}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)$ (164)
@@
$\varepsilon_{errorguess}^{achievability}\left(p_{UV},r\right)=\left(1+\frac{1}{2^{m+1}}\right)\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)-\frac{1}{2^{m+1}}\\ -\frac{2^{m}+1}{2}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)+\frac{(1+3\delta)^{n}}{2^{m+1}}\mathcal{F}\left(n,\frac{4\delta}{1+3\delta},\left\lfloor\frac{m}{2}\right\rfloor\right)$ (165)
where $r=k/n$ and m=nkm=n-k.
@@
where $\Phi\left(x\right)=1/\sqrt{2\pi}\int_{-\infty}^{x}e^{-t^{2}/2}dt$ is the cumulative distribution function of the standard normal distribution.
@@
$\mathbb{E}\left(\mathbb{I}\left(2|V|>m\right)2^{m-2|V|}\right)\\ =-2^{m}\sum_{i=0}^{n}{n\choose i}\delta^{i}4^{-i}(1-\delta)^{n-i}\mathbb{I}\left(2i>m\right)\\ =-2^{m}\left(\frac{4-3\delta}{4}\right)^{n}\sum_{i=0}^{n}{n\choose i}\left(\frac{\delta}{4-3\delta}\right)^{i}\left(\frac{4-4\delta}{4-3\delta}\right)^{n-i}\mathbb{I}\left(2i>m\right)\\ =-2^{m}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)$ (170)
@@
$\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)(J-1)2^{-m}\right)=\sum_{v}\mathbb{P}\left(V=v\right)\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)(J-1)2^{-m}\Big|V=v\right)$ (171)
@@
$\varepsilon_{errorguess}^{achievability}\left(p_{UV},r\right)\\ =\mathbb{E}\left(\mathbb{I}\left(2|V|>m\right)\left(1-2^{m-2|V|-1}-2^{-2|V|-1}\right)\right)\\ +\mathbb{E}\left(\mathbb{I}\left(2|V|\leq m\right)\left(2^{2|V|-m-1}-2^{-m-1}\right)\right)$ (173)
@@
$\mathbb{E}\left(\mathbb{I}\left(2|V|>m\right)\left(1-2^{m-2|V|-1}-2^{-2|V|-1}\right)\right)\\ =\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)-\frac{2^{m}+1}{2}\left(\frac{4-3\delta}{4}\right)^{n}\mathcal{F}\left(n,\frac{4-4\delta}{4-3\delta},n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)$ (174)
@@
$2^{-m-1}\mathbb{E}\left(\mathbb{I}\left(2|V|\leq m\right)\right)=-2^{-m-1}\mathcal{F}\left(n,\delta,\left\lfloor\frac{m}{2}\right\rfloor\right)\\ =-2^{-m-1}\left(1-\mathcal{F}\left(n,1-\delta,n-\left\lfloor\frac{m}{2}\right\rfloor-1\right)\right)$ (176)
@@ 5.5 The qubit depolarizing channel
For every $n,p$, extend the binomial CDF $\mathcal{F}\left(n,p,\cdot\right)$ from equation (163) to a piecewise linear function whose graph connects the points (1,0)(-1,0), $(0,\mathcal{F}\left(n,p,0\right))$, …, $(n,\mathcal{F}\left(n,p,n\right))$. This extension is an increasing bijection that maps [1,n][-1,n] to $[0,1]$; let $\mathcal{F}^{-1}\left(n,p,\cdot\right)$ be the inverse function.
@@
For each $i=0$, …, $n$, the events $J/4^{n}\leq\mathcal{F}\left(n,3/4,i\right)$ and $|U|\leq i$ coincide. Therefore, the CDF of $J/4^{n}$ can be extended to a piecewise linear function $0,1]\rightarrow[0,1]$ whose graph connects the points $(0,0)$, $(\mathcal{F}\left(n,3/4,0\right),\mathcal{F}\left(n,\delta,0\right))$, …, $(\mathcal{F}\left(n,3/4,n\right),\mathcal{F}\left(n,\delta,n\right))$. This can equivalently be expressed as the composition of two piecewise linear functions: $\mathcal{F}^{-1}\left(n,3/4,\cdot\right)$, whose graph connects the points (0,1)(0,-1), $(\mathcal{F}\left(n,3/4,0\right),0)$, …, $(\mathcal{F}\left(n,3/4,n\right),n)$, and $\mathcal{F}\left(n,\delta,\cdot\right)$, whose graph connects the points (1,0)(-1,0), $(0,\mathcal{F}\left(n,\delta,0\right))$, …, $(n,\mathcal{F}\left(n,\delta,n\right))$. ∎
@@
where $r=k/n$, m=nkm=n-k and $\ell=\mathcal{F}^{-1}\left(n,\frac{3}{4},\frac{2^{m}}{4^{n}}\right)$.
@@
$R_{errorguess}\left(p_{U},\epsilon\right)=1-h(\delta)-\delta\log_{2}(3)-\sqrt{\frac{\delta(1-\delta)}{n}}\Phi^{-1}\left(\epsilon\right)\log_{2}\left(\frac{\delta}{3(1-\delta)}\right)\\ +\frac{\log_{2}(n)}{2n}+O\left(\frac{1}{n}\right)$ (192)
@@
$\mathbb{P}\left(J>2^{m}\right)+\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)(J-1)2^{-m}\right)\\ =\mathbb{P}\left(J>2^{m}\right)-\frac{1}{2^{m+1}}\mathbb{P}\left(J\leq 2^{m}\right)+\frac{1}{2^{m}}\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)\left(J-\frac{1}{2}\right)\right)\\ =\left(1+\frac{1}{2^{m+1}}\right)\mathcal{F}\left(n,1-\delta,n-1-\ell\right)-\frac{1}{2^{m+1}}+\frac{1}{2^{m}}\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)\left(J-\frac{1}{2}\right)\right)$ (194)
@@
$\mathbb{E}\left(\mathbb{I}\left(4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)<J\leq 4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)\right)\left(J-\frac{1}{2}\right)\right)\\ =\mathbb{P}\left(4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)<J\leq 4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)\right)\\ *\mathbb{E}\left(J-\frac{1}{2}\Bigg|4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)<J\leq 4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)\right)\\ =\left(4^{n}\mathcal{F}\left(n,\frac{3}{4},i\right)-4^{n}\mathcal{F}\left(n,\frac{3}{4},i-1\right)\right)\\ *\left(\frac{\delta}{3}\right)^{i}(1-\delta)^{n-i}\frac{4^{n}\mathcal{F}\left(n,3/4,i-1\right)+4^{n}\mathcal{F}\left(n,3/4,i\right)}{2}\\ =\frac{(16-16\delta)^{n}}{2}\left(\frac{\delta}{3-3\delta}\right)^{i}\left(\mathcal{F}\left(n,\frac{3}{4},i\right)^{2}-\mathcal{F}\left(n,\frac{3}{4},i-1\right)^{2}\right)$ (197)
@@
$\frac{1}{2^{m}}\mathbb{E}\left(\mathbb{I}\left(J\leq 2^{m}\right)\left(J-\frac{1}{2}\right)\right)\\ =\frac{(16-16\delta)^{n}}{2^{m+1}}\left(\frac{\delta}{3-3\delta}\right)^{\left\lfloor\ell\right\rfloor+1}\mathcal{F}\left(n,\frac{3}{4},\ell\right)^{2}\\ +\frac{(16-16\delta)^{n}}{2^{m+1}}\left(\frac{3-4\delta}{3-3\delta}\right)\sum_{i=0}^{\left\lfloor\ell\right\rfloor}\left(\frac{\delta}{3-3\delta}\right)^{i}\mathcal{F}\left(n,\frac{3}{4},i\right)^{2}$ (199)
@@
$m=n(h(\delta)+\delta\log_{2}(3))+\sqrt{n\delta(1-\delta)}\Phi^{-1}\left(\epsilon\right)\log_{2}\left(\frac{\delta}{3(1-\delta)}\right)-\frac{1}{2}\log_{2}(n)+O(1)$ (202)
@@
$m=n(h(\delta)+\delta\log_{2}(3))+\sqrt{n\delta(1-\delta)}\Phi^{-1}\left(\epsilon\right)\log_{2}\left(\frac{\delta}{3(1-\delta)}\right)-\frac{1}{2}\log_{2}(n)+O(1)$ (205)
@@
${n\choose\left\lfloor\ell\right\rfloor}\left(\frac{3}{4}\right)^{\left\lfloor\ell\right\rfloor}\left(\frac{1}{4}\right)^{n-\left\lfloor\ell\right\rfloor}\leq\mathcal{F}\left(n,\frac{3}{4},\ell\right)\leq\frac{3}{2}{n\choose\left\lceil\ell\right\rceil}\left(\frac{3}{4}\right)^{\left\lceil\ell\right\rceil}\left(\frac{1}{4}\right)^{n-\left\lceil\ell\right\rceil}$ (210)
@@
and the ratio of the (i1)(i-1)-st to the $i$-th term in the sum is
$\frac{{n\choose i-1}\left(\frac{3}{4}\right)^{i-1}\left(\frac{1}{4}\right)^{n-i+1}}{{n\choose i}\left(\frac{3}{4}\right)^{i}\left(\frac{1}{4}\right)^{n-i}}=\frac{i}{3(n-i+1)}\leq\frac{2\delta}{3(1-2\delta)}\leq\frac{1}{3}$ (212)
@@
$\log_{2}\mathcal{F}\left(n,\frac{3}{4},\ell\right)=-nD\left(\frac{\ell}{n}\middle\|\frac{3}{4}\right)-\frac{1}{2}\log_{2}(n)+O(1)$ (213)
where $D\left(\cdot\middle\|\cdot\right)$ is the relative entropy.
@@
$D\left(\frac{\ell}{n}\middle\|\frac{3}{4}\right)=D\left(\delta\middle\|\frac{3}{4}\right)-\sqrt{\frac{\delta(1-\delta)}{n}}\Phi^{-1}\left(\epsilon\right)\log_{2}\left(\frac{\delta}{3(1-\delta)}\right)+O\left(\frac{1}{n}\right)$ (215)
@@
$D\left(x\middle\|y\right)=D\left(x_{0}\middle\|y\right)+(x-x_{0})\log_{2}\left(\frac{x_{0}(1-y)}{(1-x_{0})y}\right)+O\left((x-x_{0})^{2}\right)$ (216)
@@
Claim 4: (213) and (215) and the identity $2-D\left(\delta\middle\|3/4\right)=h(\delta)+\delta\log_{2}(3)$ prove (208).